# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ Abstract.
The virtual cohomological dimension of $\operatorname{Out}(F_{n})$ is given precisely by the dimension of the spine of Culler–Vogtmann Outer space. However, the dimension of the spine of untwisted Outer space for a general right-angled Artin group $A_{\Gamma}$ does not necessarily match the virtual cohomological dimension of the untwisted subgroup $U(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$. Under certain graph-theoretic conditions, we perform an equivariant deformation retraction of this spine to produce a new contractible cube complex upon which $U(A_{\Gamma})$ acts properly and cocompactly. Furthermore, we give conditions for when the dimension of this complex realises the virtual cohomological dimension of $U(A_{\Gamma})$.
@@ 1.1. Overview of results
For any right-angled Artin group $A_{\Gamma}$, Charney–Stambaugh–Vogtmann [CSV17] constructed an untwisted Outer space $\Sigma_{\Gamma}$, which is a contractible complex with a proper $U(A_{\Gamma})$-action. This is a generalisation of the influential Culler–Vogtmann Outer space $CV_{n}$, introduced in 1986 [CV86]. Untwisted Outer space has a spine $K_{\Gamma}$—a deformation retract of $\Sigma_{\Gamma}$ which naturally has the structure of a cube complex and which has a proper and cocompact $U(A_{\Gamma})$-action. This implies that $\dim\left(K_{\Gamma}\right)$ is an upper bound for the virtual cohomological dimension (vcd) of $U(A_{\Gamma})$.
In many cases, equality holds. Millard and Vogtmann [MV19] show that the principal rank (see Definition 3.12) of $\Gamma$ is a lower bound for $\textsc{vcd}(U(A_{\Gamma}))$. However, there exist simplicial graphs $\Gamma$ for which the principal rank is strictly less than $\dim(K_{\Gamma})$. For example, the 2-rake graph $T_{2}$ shown in Figure 1 has principal rank equal to 5 but $\dim(K_{T_{2}})=6$.
$u$$v$$a_{1}$$a_{2}$$b_{1}$$b_{2}$
@@
In [MV19], Millard and Vogtmann give a sufficient condition for the principal rank to be equal to $\dim(K_{\Gamma})$ (and hence for $\textsc{vcd}(U(A_{\Gamma}))=\dim(K_{\Gamma})$). This graph-theoretic condition roughly says that all the vertices of $\Gamma$ which ‘dominate’ another vertex $u$ lie in the same component of $\Gamma\setminus\operatorname{st}({u})$ (here the star of $u$$\operatorname{st}({u})$, is the full subgraph on $u$ and all its neighbours; by ‘dominate’ we refer to the relation $<_{\circ}$—see §2.2). One may wonder whether this hypothesis can be weakened to give a sufficient condition guaranteeing that the gap between $\textsc{vcd}(U(A_{\Gamma}))$ and $\dim\left(K_{\Gamma}\right)$ is at most $k\geq 1$. Arguably the most natural guess for such a condition is the following condition, which we denote $P(k)$: the vertices ‘dominating’ $u$ must be divided between at most $(k+1)$ components of $\Gamma\setminus\operatorname{st}({u})$. Indeed, our prototypical examples of graphs which satisfy $P(k)$ (see §6.1) indeed have $\dim\left(K_{\Gamma}\right)-\textsc{vcd}(U(A_{\Gamma}))$ equal to exactly $k$. However, we dispel this speculation by applying (the proof of) Theorem A to $\Delta$, to obtain:
@@ 1.2. Strategy of proofs
We introduce the idea of 1- and 2-hugged $\Gamma$-partitions (see Definition 4.2), which roughly captures a notion of ’redundancy’ among the cubes of $K_{\Gamma}$. Most of the work of this paper goes into the following pair of statements.
Theorem D 4, 5.4).
@@
    Consider the subcomplex $\widehat{K_{\Gamma}}$ of $K_{\Gamma}$ consisting only of cubes which are not redundant. If $\Gamma$ is spiky, then $\smash{\widehat{K_{\Gamma}}}$ can be obtained as an equivariant deformation retraction of $K_{\Gamma}$.
@@
The $d$-rake graph $T_{d}$ (see Figure 12) has principal rank $13d-$ 1. The dimension of the spine $K_{T_{d}}$ is $24d-$ 2.
The rake graphs are both spiky and barbed, so Theorem A applies. Hence, we find that the virtual cohomological dimension $\textsc{vcd}(U(A_{T_{d}}))$ is equal to $13d-$ 1, which gives Corollary B. Certain derivatives of ‘rake’ graphs also exhibit this behaviour. Similarly, we use a good understanding of the possible $\Delta$-Whitehead partitions to obtain Corollary C.
@@ 1.3. Context
The untwisted subgroup $U(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$ is obtained by excluding certain automorphisms called twists. Twists are automorphisms which multiply a generator by another generator with which it commutes, so the untwisted subgroup is most closely related to the ‘free group end’ of the spectrum of right-angled Artin groups—where there are no commuting generators. In particular, $U(F_{n})$ is the entire group of outer automorphisms $\operatorname{Out}(F_{n})$. The abelianisation $A_{\Gamma}\to\mathbb{Z}^{n}$ induces a map $\operatorname{Out}(A_{\Gamma})\to\operatorname{GL}_{n}(\mathbb{Z})$ whose kernel is always contained in $U(A_{\Gamma})$ (this kernel was studied in more detail by Toinet, [Toi13]).
The spine for the free group $F_{n}$ has dimension $32n-$ 3, and it is easy to find a free abelian subgroup of $\operatorname{Out}(F_{n})$ of rank $32n-$ 3. Hence, the virtual cohomological dimension of $\operatorname{Out}(F_{n})$ is precisely the dimension of the spine (this was observed in [CV86]). We refer to Vogtmann’s surveys [Vog02], [Vog18], and [Vog] for detailed descriptions of Culler–Vogtmann Outer space and its spine, as well as their applications.
@@
Building on the untwisted Outer space constructed in [CSV17], Bregman–Charney–Vogtmann [BCV23] constructed a full Outer space $\mathcal{O}_{\Gamma}$ – a contractible space with a proper $\operatorname{Out}(A_{\Gamma})$-action. This has already been used, for example, to study finite subgroups of $U^{0}(A_{\Gamma})$, a finite index subgroup of $U(A_{\Gamma})$: see [BCV24].
@@
Regarding the first of these possibilities, it is proved in [CV11] that for many graphs, $U(A_{\Gamma})$ cannot contain a torsion-free non-abelian solvable subgroup. However, there still exist graphs for which we cannot disregard such a possibility (compare with $\textsc{vcd}(\operatorname{GL}_{n}(\mathbb{Z}))$, which is realised as the Hirsch rank of a (non-abelian) polycyclic subgroup). As yet, we know of no example where $\textsc{vcd}(U(A_{\Gamma}))$ is not realised as the rank of a (natural) free abelian subgroup. (Millard–Vogtmann’s proof [MV19] that $M(L)\leq\textsc{vcd}(U(A_{\Gamma}))$ proceeds by finding a free abelian subgroup of rank $M(L)$, generated in a rather natural way.)
@@ 1.4. Further directions
There are certain obvious questions regarding the approach taken here. It is not known to what extent the conditions spiky and barbed are simply artefacts of the proof. In §4.3 and §7, we discuss examples of graphs that violate spikiness in different ways. For one of these examples (Figure 13), our results can nevertheless be applied with no modification. However, for another (Eg. 4.9), the retraction process cannot be executed in its current form. This requires further investigation—for example, the notion of redundancy used in this paper can be tweaked to be applied to the graph in Eg. 4.9, but this comes at a cost of a less streamlined definition and a less intuitive (if slightly weaker) version of our ‘spikiness’ condition.
There are also natural questions to be asked about our resultant complex $\widehat{K_{\Gamma}}$. For example, is $\dim(\widehat{K_{\Gamma}})$ always equal to $\textsc{vcd}(U(A_{\Gamma}))$? At present, we know of no example to the contrary. In another direction, Bridson–Vogtmann [BV01] proved that $\operatorname{Out}(F_{n})$ is the group of simplicial automorphisms of the corresponding spine (for $n\geq 3$). We may ask the same question here: when is the group of simplicial automorphisms of $K_{\Gamma}$ isomorphic to $U(A_{\Gamma})$? What about the group of automorphisms of $\widehat{K_{\Gamma}}$? In other words: when, if ever, is $K_{\Gamma}$ or $\widehat{K_{\Gamma}}$ an accurate model for $U(A_{\Gamma})$?
Finally, we remark that there already exists an algorithm, due to Day–Sale–Wade [DSW19], which computes the virtual cohomological dimension of $\operatorname{Out}(A_{\Gamma})$ and $U(A_{\Gamma})$ for any graph $\Gamma$. This algorithm is difficult to apply in practice and does not provide a complex whose dimension realises the virtual cohomological dimension. One interesting further research direction would be to compare and contrast these two approaches for computing $\textsc{vcd}(U(A_{\Gamma}))$. For example, is there any transparent way of translating the value of $\textsc{vcd}(U(A_{\Gamma}))$ calculated in [DSW19] into the size of a compatible set of $\Gamma$-Whitehead partitions? If so, does this set resemble the top-dimensional cubes in $\widehat{K_{\Gamma}}$ in any way?
@@
Is $\textsc{vcd}(U(A_{\Gamma}))$ equal to the principal rank of $\Gamma$?
The results in this paper give a positive answer to this question for spiky, barbed $\Gamma$. Strengthening the approach taken here would therefore be one way of looking for a positive answer in general. On the other hand, one may search for counterexamples by applying Day–Sale–Wade’s algorithm to graphs which fail to be both spiky and barbed.
@@
In §2, we establish notation and gather the necessary preliminaries. §3 introduces the spine of untwisted Outer space and discusses the work of [MV19] regarding the principal rank. In §4, we define ‘spikiness’ and detail the retraction process which obtains $\widehat{K_{\Gamma}}$ from $K_{\Gamma}$, proving Theorem D (i). §5 is dedicated to proving Theorem D (ii) and deducing Theorem A. Finally, in §6 and §7, we present detailed examples and computations, deducing Corollaries B & C.
@@ Acknowledgements
I would also like to thank my doctoral supervisors, Rachael Boyd and Tara Brendle, for their guidance, patience, and trust, as well as helpful discussions. Thanks are also due to Corey Bregman for taking the time to read through an earlier draft of this paper and for his helpful comments.
@@ 2.1. Right-angled Artin groups
Let $\Gamma$ be a finite simplicial graph (that is, it has no loops and no multi-edges) whose vertices are labelled. Write $V(\Gamma)$ and $E(\Gamma)$ for its vertex and edge sets, respectively. The associated right-angled Artin group (RAAG), denoted $A_{\Gamma}$, is defined by the presentation
@@
$\Gamma$ is the defining graph for $A_{\Gamma}$. We will blur the distinction between a vertex of $\Gamma$ and the corresponding generator of $A_{\Gamma}$, and will communicate the data of a RAAG simply by its defining graph.
Every Artin group has an associated cell complex called the Salvetti complex [Sal87]. For a RAAG $A_{\Gamma}$, the Salvetti complex $\mathbb{S}_{\Gamma}$ is a cube complex with a particularly simple description, as follows. There is one 0-cell, to which is attached a directed 1-cell for each generator. Then for each edge $\{a,b\}$ in $\Gamma$, we glue in a 2-torus along the corresponding commutator $aba^{-1}b^{-1}$. Continuing inductively, for each $k$-clique in $\Gamma$ (which corresponds to a set of $k$ mutually commuting generators), we glue in a $k$-torus whose constituent $(k-1)$-tori correspond to the subcliques of our $k$-clique which have $k-1$ vertices.
For example, the Salvetti complex of the free group of rank $n$ is simply the $n$-petalled rose graph, while the Salvetti complex of $\mathbb{Z}^{n}$ is an $n$-torus. The Salvetti complex of a RAAG associated to a triangle-free graph is the presentation complex of the RAAG. Note that $\pi_{1}(\mathbb{S}_{\Gamma})\cong A_{\Gamma}$. Charney and Davis proved [CD96] that $\mathbb{S}_{\Gamma}$ is a $K(A_{\Gamma},1)$ space, resolving the $K(\pi,1)$ conjecture for RAAGs; it follows that RAAGs are torsion-free and biautomatic (as shown by Niblo and Reeves [NR97]), for instance.
@@ 2.2. Graph-theoretic preliminaries
We define a relation $\leq$ on the vertices of $\Gamma$ by saying that $v\leq w$ if $\operatorname{lk}({v})$ is contained in $\operatorname{st}({w})$. We may define an equivalence relation $\sim$ on $V(\Gamma)$ by saying that $v\sim w$ if both $v\leq w$ and $w\leq v$. Then $\leq$ is a partial order on the set of equivalence classes. We will write $[v]$ for the equivalence class of $v$.
@@
All maximal vertices are principal. As observed in [MV19], not all principal vertices are maximal. For example, in the triangle with leaves at two of its vertices, the third vertex of the triangle is principal but not maximal.
@@ 2.3. Automorphisms of RAAGs
For any $\Gamma$$\operatorname{Aut}(A_{\Gamma})$ is generated by the following set of elementary transformations.
@@
    • folds if $v,w$ are not adjacent. In particular, $v\mapsto vw$ is a right fold, while $v\mapsto wv$ is a left fold.
@@
    Partial conjugations: for a vertex $v$, let $C$ be a connected component of $\Gamma\setminus\operatorname{st}({v})$. The map which sends every generator $c\in C$ to the conjugation $vcv^{-1}$ and fixes all other generators is called a partial conjugation (Laurence and Servatius refer to these as locally inner automorphisms).

### v2

@@ Abstract.
The virtual cohomological dimension of $\operatorname{Out}(F_{n})$ is given precisely by the dimension of the spine of Culler–Vogtmann Outer space. However, the dimension of the spine of untwisted Outer space for a general right-angled Artin group $A_{\Gamma}$ does not necessarily match the virtual cohomological dimension $\textsc{vcd}(U(A_{\Gamma}))$ of the untwisted subgroup $U(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$. Under certain graph-theoretic conditions, we perform an equivariant deformation retraction of this spine to produce a new contractible cube complex upon which $U(A_{\Gamma})$ acts properly and cocompactly. Furthermore, we give conditions for when the dimension of this complex realises the virtual cohomological dimension of $U(A_{\Gamma})$. We finish with two applications of our construction; in particular we show that the difference between the dimension of the untwisted spine and $\textsc{vcd}(U(A_{\Gamma}))$ can be arbitrarily large.
@@ 1.1. Overview of results
For any right-angled Artin group $A_{\Gamma}$, Charney–Stambaugh–Vogtmann [CSV17] constructed an untwisted Outer space $\mathcal{O}_{\Gamma}^{U}$, which is a contractible complex with a proper $U(A_{\Gamma})$-action. This is a generalisation of the influential Culler–Vogtmann Outer space $CV_{n}$, introduced in 1986 [CV86]. Untwisted Outer space has a spine $K_{\Gamma}$ – a deformation retract of $\mathcal{O}_{\Gamma}^{U}$ which naturally has the structure of a cube complex and which has a proper and cocompact $U(A_{\Gamma})$-action. This implies that $\dim\left(K_{\Gamma}\right)$ is an upper bound for the virtual cohomological dimension (vcd) of $U(A_{\Gamma})$.
In many cases, equality holds. Millard and Vogtmann [MV21] show that the principal rank (see Definition 3.12) of $\Gamma$ is a lower bound for $\textsc{vcd}(U(A_{\Gamma}))$. However, there exist simplicial graphs $\Gamma$ for which the principal rank is strictly less than $\dim(K_{\Gamma})$. For example, the 2-rake graph $T_{2}$ shown in Figure 1 has principal rank equal to 5 but $\dim(K_{T_{2}})=6$.
[…]
@@
[…]
@@ 1.2. Strategy of proofs
We introduce the idea of 1- and 2-hugged $\Gamma$-partitions (see Definition 4.2), which roughly captures a notion of ‘redundancy’ among the cubes of $K_{\Gamma}$. Most of the work of this paper goes into the following pair of statements.
Theorem D 4, 5.5).
@@
    Consider the subcomplex $\widehat{K_{\Gamma}}$ of $K_{\Gamma}$ consisting only of cubes which are not redundant. If $\Gamma$ is spiky, then $\smash{\widehat{K_{\Gamma}}}$ can be obtained as a $U(A_{\Gamma})$-equivariant deformation retraction of $K_{\Gamma}$.
@@
The $d$-rake graph $T_{d}$ (see Figure 12) has principal rank 3d13d-1. The dimension of the spine $K_{T_{d}}$ is 4d24d-2.
The rake graphs are both spiky and barbed, so Theorem A applies. Hence, we find that the virtual cohomological dimension $\textsc{vcd}(U(A_{T_{d}}))$ is equal to 3d13d-1, which gives Corollary B. Certain derivatives of ‘rake’ graphs also exhibit this behaviour. Similarly, we use a good understanding of the possible $\Delta$-Whitehead partitions to obtain Corollary C.
@@ 1.3. Context
The untwisted subgroup $U(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$ is obtained by excluding twists. Twists are automorphisms which multiply a generator by another generator with which it commutes, so the untwisted subgroup is most closely related to the ‘free group end’ of the spectrum of right-angled Artin groups – where there are no commuting generators. In particular, $U(F_{n})$ is the entire group of outer automorphisms $\operatorname{Out}(F_{n})$. The abelianisation $A_{\Gamma}\to\mathbb{Z}^{n}$ induces a map $\operatorname{Out}(A_{\Gamma})\to\operatorname{GL}_{n}(\mathbb{Z})$ whose kernel is always contained in $U(A_{\Gamma})$ (this kernel was studied in more detail by Toinet, [TOI13]).
The spine for the free group $F_{n}$ has dimension 2n32n-3, and it is easy to find a free abelian subgroup of $\operatorname{Out}(F_{n})$ of rank 2n32n-3. Hence, the virtual cohomological dimension of $\operatorname{Out}(F_{n})$ is precisely the dimension of the spine (this was observed in [CV86]). We refer to Vogtmann’s surveys [VOG02], [VOG18], and [CHV25] for detailed descriptions of Culler–Vogtmann Outer space and its spine, as well as their applications.
@@
Building on the untwisted Outer space constructed in [CSV17], Bregman–Charney–Vogtmann [BCV23] constructed a full Outer space $\mathcal{O}_{\Gamma}$ – a contractible space with a proper $\operatorname{Out}(A_{\Gamma})$-action.
@@
Regarding the first of these possibilities, it is proved in [CV11] that for many graphs, $U(A_{\Gamma})$ cannot contain a torsion-free non-abelian solvable subgroup. However, there still exist graphs for which we cannot disregard such a possibility (compare with $\textsc{vcd}(\operatorname{GL}_{n}(\mathbb{Z}))$, which is realised as the Hirsch rank of a (non-abelian) polycyclic subgroup (see [BIE81, Theorem 7.10 & §7.4])). As yet, we know of no example where $\textsc{vcd}(U(A_{\Gamma}))$ is not realised as the rank of a (natural, easily found) free abelian subgroup (Millard–Vogtmann’s proof [MV21] that $M(L)\leq\textsc{vcd}(U(A_{\Gamma}))$ proceeds by finding a free abelian subgroup of rank $M(L)$, generated in a reasonably anticapable way.)
@@ 1.4. Further directions
There are certain obvious questions regarding the approach taken here. It is not known to what extent the conditions spiky and barbed are simply artefacts of the proof. In §4.3 and §7, we discuss examples of graphs that violate spikiness in different ways. For one of these examples (Figure 13), our results can nevertheless be applied with no modification. However, for another (Eg. 4.10), the retraction process cannot be executed in its current form. This requires further investigation – for example, the notion of redundancy used in this paper can be tweaked to be applied to the graph in Eg. 4.10, but this comes at the cost of a less streamlined definition and a less intuitive (if slightly weaker) version of our ‘spikiness’ condition.
[…]
Finally, we remark that there already exists an algorithm, due to Day–Sale–Wade [DSW21], which computes the virtual cohomological dimension of $\operatorname{Out}(A_{\Gamma})$ and $U(A_{\Gamma})$ for any graph $\Gamma$. This algorithm is difficult to apply in practice and does not provide a complex whose dimension realises the virtual cohomological dimension. One interesting further research direction would be to compare and contrast these two approaches for computing $\textsc{vcd}(U(A_{\Gamma}))$. For example, is there any transparent way of translating the value of $\textsc{vcd}(U(A_{\Gamma}))$ calculated in [DSW21] into the size of a compatible set of $\Gamma$-Whitehead partitions? If so, does this set resemble the top-dimensional cubes in $\widehat{K_{\Gamma}}$ in any way?
@@
Is $\textsc{vcd}(U(A_{\Gamma}))$ always equal to the principal rank of $\Gamma$?
The results in this paper give a positive answer to this question for spiky, barbed $\Gamma$. Strengthening the approach taken here would therefore be one way of looking for a positive answer in general. On the other hand, one may search for counterexamples by applying Day–Sale–Wade’s algorithm to graphs which fail to be simultaneously spiky and barbed.
@@
In §2, we establish notation and gather the necessary preliminaries. §3 introduces the spine of untwisted Outer space and discusses the work of [MV21] regarding the principal rank. In §4, we define ‘spikiness’ and detail the retraction process which obtains $\widehat{K_{\Gamma}}$ from $K_{\Gamma}$, proving Theorem D (i). §5 is dedicated to proving Theorem D (ii) and deducing Theorem A. Finally, in §6 and §7, we present detailed examples and computations, deducing Corollaries B & C.
@@ Acknowledgements
I would also like to thank my doctoral supervisors, Rachael Boyd and Tara Brendle, for their guidance, patience, and trust, as well as helpful discussions. Thanks are also due to Corey Bregman for taking the time to read through an earlier draft of this paper and for his helpful comments. Finally, I would like to express my gratitude to the referee for their careful reading and many helpful suggestions for improvement.
@@ 2.1. Right-angled Artin groups
Let $\Gamma$ be a finite simplicial graph (that is, it has no single-edge loops and no multi-edges) whose vertices are labelled. Write $V(\Gamma)$ and $E(\Gamma)$ for its vertex and edge sets, respectively. The associated right-angled Artin group (RAAG), denoted $A_{\Gamma}$, is defined by the presentation
@@
Say that $\Gamma$ is the defining graph for $A_{\Gamma}$. We will blur the distinction between a vertex of $\Gamma$ and the corresponding generator of $A_{\Gamma}$, and will communicate the data of a RAAG simply by its defining graph.
Every Artin group has an associated cell complex called the Salvetti complex [SAL87]. For a RAAG $A_{\Gamma}$, the Salvetti complex $\mathbb{S}_{\Gamma}$ is a cube complex with a particularly simple description, as follows. There is one 0-cell, to which is attached a directed 1-cell for each generator. Then for each edge $\{a,b\}$ in $\Gamma$, we glue in a 2-torus along the corresponding commutator $aba^{-1}b^{-1}$. Continuing inductively, for each $k$-clique in $\Gamma$ (which corresponds to a set of $k$ mutually commuting generators), we glue in a $k$-torus whose constituent (k1)(k-1)-tori correspond to the subcliques of our $k$-clique which have k1k-1 vertices.
For example, the Salvetti complex of the free group of rank $n$ is simply the $n$-petalled rose graph, while the Salvetti complex of $\mathbb{Z}^{n}$ is an $n$-torus. The Salvetti complex of a RAAG associated to a triangle-free graph is the presentation complex of the RAAG. Note that $\pi_{1}(\mathbb{S}_{\Gamma})\cong A_{\Gamma}$. Charney and Davis proved [CD95] that $\mathbb{S}_{\Gamma}$ is a $K(A_{\Gamma},1)$ space, resolving the $K(\pi,1)$ conjecture for RAAGs; it follows that RAAGs are torsion-free and biautomatic (as shown by Niblo and Reeves [NR98]), for instance.
@@ 2.2. Graph-theoretic preliminaries
Define a relation $\leq$ on the vertices of $\Gamma$ by saying that $v\leq w$ if $\operatorname{lk}({v})$ is contained in $\operatorname{st}({w})$. We define an equivalence relation $\sim$ on $V(\Gamma)$ by saying that $v\sim w$ if both $v\leq w$ and $w\leq v$. Then $\leq$ is a partial order on the set of equivalence classes. We write $[v]$ for the equivalence class of $v$.
@@
All maximal vertices are principal. However, as observed in [MV21], not all principal vertices are maximal. For example, in the triangle with leaves at two of its vertices, the third vertex of the triangle is principal but not maximal (see Figure 3).
@@ 2.3. Automorphisms of RAAGs
[…]
@@
    • folds, if $v,w$ are not adjacent. In particular, $v\mapsto vw$ is a right fold, while $v\mapsto wv$ is a left fold.
@@
    Partial conjugations: for a vertex $v$, let $C$ be a connected component of $\Gamma\setminus\operatorname{st}({v})$. The map which sends every generator $c\in C$ to the conjugate $vcv^{-1}$ and fixes all other generators is called a partial conjugation (Laurence and Servatius refer to these as locally inner automorphisms).
When talking about generators of $\operatorname{Out}(A_{\Gamma})$ and its subgroups, we will use the same terminology: e.g., by a ‘partial conjugation’ in $\operatorname{Out}(A_{\Gamma})$ we mean the image of a partial conjugation under the quotient $\operatorname{Aut}(A_{\Gamma})\twoheadrightarrow\operatorname{Out}(A_{\Gamma})$.